3.) For a 0.12 M solution of HCl (aq), calculate the following: a. [H3O+] b. pH c. pOH d. [OH-]
Answer #1

Given Conc. of HCl = 0.12 M

So;

a) [H3O+] = 0.12

b) pH = - log[H+] = - log [0.12] = 0.92

c) pH + pOH = 14; pOH = 14 - pH = 14 - 0.92 = 13.08

d) [OH-] = 10-pOH = 10-13.08 = 8.3 x 10-14